Sol Summary – April 20th

Crew 245 Sol Summary Report 20-04-2022
SOL: 10
Summary Title: A good sol
Author’s name: Paolo Guardabasso

Mission status: Operations proceed nominally, with some more focus on the journalist’s work filming the station.

Sol Activity Summary: This morning the cortisol stress level test experiment EVA was completed, and in the afternoon we dedicated time to experiments and return logistics.

Look Ahead Plan: Tomorrow will be our last day of sim. We plan to spend the day doing inventory for shipment, work on the final research report and perhaps perform our two last EVAs, depending on the weather.

Anomalies in work: None, the suit problem apparently had a very trivial solution, linked to a misplacement of the chargers’ multiple plugs (see operations report for details).

Weather: Very beautiful day, that allowed swift EVAs and comfortable temperature in the afternoon. Temperature 6 (low) and 25 (high) C, Wind speed 19 (avg) and 24 (max) kph

Crew Physical Status: All nominal.

EVA: Two consecutive EVAs have been performed. They represent the second part of two parts of the stress level experiment, where 2 crew members will have to explore and navigate an area with and without a digital map previously generated with a drone.

Reports to be filled:
– HSO report
– Operation report
– Journalist report
– Some pictures + picture of the day
– EVA request (2)
– EVA report (2)

Support Requested: None at the moment.

Copyright © The Mars Society. All rights reserved. | Main Site